Suppose you are the money manager of a $3.74 million investment fund. The fund consists of four stocks with the following investments and betas: Stock Investment Beta A $ 320,000 1.50 B 800,000 (0.50 ) C 920,000 1.25 D 1,700,000 0.75 If the market's required rate of return is 13% and the risk-free rate is 6%, what is the fund's required rate of return? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1
Total Fund value = Value of A + Value of B + Value of C + Value of D
=320000+800000+920000+1700000
=3740000
Weight of A = Value of A/Total Fund Value
= 320000/3740000
=0.0856
Weight of B = Value of B/Total Fund Value
= 800000/3740000
=0.2139
Weight of C = Value of C/Total Fund Value
= 920000/3740000
=0.246
Weight of D = Value of D/Total Fund Value
= 1700000/3740000
=0.4545
Beta of Fund = Weight of A*Beta of A+Weight of B*Beta of B+Weight of C*Beta of C+Weight of D*Beta of D
Beta of Fund = 1.5*0.0856+-0.5*0.2139+1.25*0.246+0.75*0.4545
Beta of Fund = 0.669786
As per CAPM
expected return = risk-free rate + beta * (expected return on the market - risk-free rate)
Expected return% = 6 + 0.669786 * (12 - 6)
Expected return% = 10.02